OpenStudy (anonymous):

A bag contains 6 cherry, 3 orange, and 2 lemon candies. You reach in and take 3 pieces of candy at random. The probability that you have 1 cherry candy and 2 lemon candies is

OpenStudy (anonymous):

To calculate the gross probability of drawin those 3 candies you take the ratio of those individual probabilies and multiply them together. Since you are taking them all out at the same time, you'll keep the denominator of the ratios the same (ie. 6/11 pieces of candy for cherry) since it is the total number of candies. so, if I remember correctly, you'll be multiplying 6/11*2/11*2/11. That should give you the right probability.

OpenStudy (anonymous):

Thanks let me try working it

OpenStudy (anonymous):

6/11*2/11*2/11=12/1331?

OpenStudy (anonymous):

So this is for my finals is that what I will enter for-> The standard quota for the College of Sciences is

OpenStudy (anonymous):

or 0.00901577761081893?

OpenStudy (anonymous):

6*2*2= ?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

24/11

OpenStudy (anonymous):

First off, it would stay in the large post-multiplication denominator. so 24/1331. As for your answers form, I would recommend either a fraction in simplest terms or a percentage. Decimals are not typically used for a probability answer, but it's really easy to convert them into a percentage.

OpenStudy (anonymous):

so is my anwser right? 24/1331 or 24/11

OpenStudy (ganpat):

Number of ways 3 candies can be picked = 11C3 = (11 * 10 * 9 * 8!) / (3 * 2 * 8!) = 165 Number of ways for 1 cherry = 6 C1 = 6 Number of ways for two candies = 2 C2 = 1 So the answer will be = (6 *2)/ 165 = 12 / 165..

OpenStudy (anonymous):

I was way off :-/

OpenStudy (anonymous):

Thank u so much!!!...

OpenStudy (ganpat):

@bl4ckj4ck777 : can you check ma reply ??